# The Sound of Silence in Social Networks

### v1

@@ Abstract
We generalize the classic multi-agent DeGroot model for opinion dynamics to incorporate the Spiral of Silence theory from political science. This theory states that individuals may withhold their opinions when they perceive them to be in the minority. As in the DeGroot model, a community of agents is represented as a weighted directed graph whose edges indicate how much agents influence one another. However, agents whose current opinions are in the minority become silent (i.e., they do not express their opinion). Two models for opinion update are then introduced. In the memoryless opinion model ($\mbox{SOM}^{-}$), agents update their opinion by taking the weighted average of their non-silent neighbors’ opinions. In the memory based opinion model ($\mbox{SOM}^{+}$), agents update their opinions by taking the weighted average of the opinions of all their neighbors, but for silent neighbors, their most recent opinion is considered.
We show that for $\mbox{SOM}^{-}$ convergence to consensus is guaranteed for clique graphs but, unlike for the classic DeGroot, not guaranteed for strongly-connected aperiodic graphs. In contrast, we show that for $\mbox{SOM}^{+}$ convergence to consensus is not guaranteed even for clique graphs. We showcase our models through simulations offering experimental insights that align with key aspects of the Spiral of Silence theory. These findings reveal the impact of silence dynamics on opinion formation and highlight the limitations of consensus in more nuanced social models.
Keywords:
Opinion Dynamics, Spiral of Silence, DeGroot Model, Social Networks, Consensus, Agent-Based Modeling, Social Influence
@@ 1 Introduction
Social networks have played a significant role in shaping users’ opinions, often influencing democratic processes and contributing to social polarization. Broadly, the dynamics of opinion formation in social networks involve users expressing their opinions, being exposed to the opinions of others, and potentially adapting their own views based on these interactions. Modeling these dynamics enables us to glean insights into how opinions form and spread within social networks.
The DeGroot model [12] is one of the most prominent formalisms for opinion formation and consensus-building in social networks. In this model, a social network is represented as a weighted directed graph, where edges denote the degree to which individuals (referred to as agents) influence one another. Each agent holds an opinion, expressed as a value in $[0,1]$, indicating their level of agreement with an underlying proposition (e.g., “AI is a threat to humanity"). Agents repeatedly update their opinions by taking the weighted average of their opinion differences with those who influence them (i.e, their neighbours or contacts). There is empirical evidence validating the opinion formation through averaging of the model in controlled sociological experiments (e.g., [8]).
Consensus, i.e., convergence to a common opinion, is a central property in models of social learning and opinion formation [16]. In fact, difficulties in achieving consensus are a sign of a polarized society. A fundamental result in the DeGroot model shows that agents converge to consensus if the influence graph is strongly connected and aperiodic. The DeGroot model is recognized for its mathematical simplicity, derived from its associations with matrix powers and Markov chains. It continues to be focus of research for constructing frameworks for understanding opinion formation dynamics in social networks (e.g,. [16, 5, 15, 20, 2, 11, 10, 27, 26, 13, 9, 4, 3, 22]).
Nevertheless, the DeGroot model makes an assumption that could be overly constraining within social network contexts. It assumes that all agents express their opinions at each time unit. This assumption, which renders the model tractable, may hold in some controlled scenarios as participants are often encouraged to express their views freely and consistently. However, in many real-world situations, some individuals may choose not to express their opinions due to personal choice or social pressure.
@@
In this paper, we generalize the classic DeGroot model into a framework where agents may choose to remain silent at a given time unit in line with the Spiral of Silence theory. We consider two possibilities, leading to the two models described below.
@@
Memoryless framework $\mbox{SOM}^{-}$. In this model, silent agents are excluded from the opinion updates of the agents they would typically influence. Additionally, agents become silent at a given time (thus withholding their opinions) if their views do not align with the majority of their non-silent contacts. This framework is called the memoryless silence opinion model ($\mbox{SOM}^{-}$), as the previous opinions of silent agents are not retained. This corresponds to a social scenario in which opinions (expressed as messages, posts, etc.) are removed once they have been accessed.
@@
Notice that ignoring silent agents at a given time unit amounts to removing certain edges from the underlying influence graph at that time. Thus, a fundamental distinction from the DeGroot model is that $\mbox{SOM}^{-}$ exhibits dynamic influence; i.e., edges may disappear and reappear during opinion evolution.
@@
Memory-based framework $\mbox{SOM}^{+}$. In this model, agents choose to be silent if their opinion does not align with the most recent public opinions of the majority of their contacts. In addition, when their current opinion is unknown, their most recent public opinion is taken into account in the update. Furthermore, silent agents are not excluded from the opinion updates of the agents they influence. This framework is called the memory-based silence opinion model ($\mbox{SOM}^{-}$), as some previous opinions are retained.
@@
A property that distinguishes $\mbox{SOM}^{+}$ from the DeGroot model (and $\mbox{SOM}^{-}$) is that the latter is a Markovian process: The next state depends on the current state but not the past states. Thus, $\mbox{SOM}^{+}$ is a history-dependent model but with limited memory; only the most recent public opinions are retained. We will show that this limited memory will have an impact in opinion formation and consensus.
Contributions. The main contributions in this paper are the following: (1) We generalize the DeGroot model to account for relevant social phenomena described by the theory of the Spiral of Silence. To the best of our knowledge, this is the first extension of the DeGroot model that incorporates this theory. (2) We show that in $\mbox{SOM}^{-}$, convergence to consensus is guaranteed in clique graphs (i.e., graphs where each pair of agents influences each other) with more than two agents. This intuitively means that in fully connected communities, consensus can be achieved under the Spiral of Silence assumptions, even if silent agents are excluded from the update process. (3) We show that in $\mbox{SOM}^{-}$, unlike in the DeGroot model, convergence to consensus is not guaranteed for strongly connected aperiodic graphs. (4) We demonstrate that in $\mbox{SOM}^{+}$, convergence to consensus is not guaranteed, even for cliques. This shows that, under the Spiral of Silence Theory, even the limited memory in $\mbox{SOM}^{+}$ can significantly increase the complexity of opinion dynamics, making consensus more difficult to achieve. Finally, (5) we also demonstrate our model with examples and computer simulations that provide insights into opinion formation under the Spiral of Silence. We provide examples that align with assertions of the theory of Spiral of Silence: In particular, reinforment of dominants view in social networks. The open code for these simulations can be provided upon request.
@@
All in all, this paper highlights the impact of silence dynamics and memory on opinion formation and highlight the limitations of consensus in more nuanced social models.
@@
The paper is organized as follows: The new silence opinion models are introduced in Section 2. The study of consensus for these models is presented in Section 3. Some case studies emerging from our models and highlighting the spiral of silence effect are presented in Section 4. The concluding remarks are given in Section 5.
2 The Model
@@
In the DeGroot model [12], each agent updates their opinion by taking the weighted average of the opinions of those who influence them. This model, however, does not account for the social phenomenon known as the Spiral of Silence [21], where some agents may choose to become or remain silent if their opinion does not align with the majority. As a result, their current opinion may not influence their contacts.
@@
In this section, we generalize the DeGroot model to take into account the Spiral of Silence. If an agent $j$ decides to be silent, there are at least two natural options when updating the opinions of the agents having $j$ as a contact: (1) agent $j$ is simply ignored in the update since their current opinion is unknown (or not public), or (2) the most recent opinion when $j$ was not silent is taken into account in the update. The former corresponds to a scenario where, for privacy purposes, opinions (messages) are removed once they have been accessed. The latter represents a typical scenario in social networks where previous opinions are kept and thus continue to influence others despite the agent’s current silence.
@@
The above options lead us to the two generalizations of the DeGroot model studied in this paper: the memoryless silence opinion model $\mbox{SOM}^{-}$, where previous opinions are forgotten, and the memory-based (or history-dependent) silence opinion model $\mbox{SOM}^{+}$, where previous opinions are remembered. In both generalizations, agents become silent by a majority rule appropriate for each case.
In the next sections, we introduce the elements of the models.
@@ 2.1 The Influence Graph
An ($n$-agent) influence graph is a weighted directed graph $G=(A,E,I)$, where $A=\{1,\ldots,n\}$, $E\subseteq A\times A$, and $I:A\times A\rightarrow[0,1]$ a weight function s.t. such that $I(i,j)=0{\rm\ iff\ }(i,j)\not\in E$ and for each $i\in A$, $\sum_{j\in N(i)\cup\{i\}}I(j,i)=1$ where $N_{i}=\{j\in A\setminus\{i\}:(j,i)\in E\}.$
The vertices in $A$ represent $n$ agents of a given community or network. The set of edges $E\subseteq A\times A$ represents the (direct) influence relation between these agents; i.e., $(i,j)\in E$ means that agent $i$ (directly) influences agent $j$. The value $I(i,j)$, for simplicity written $I_{ij}$, denotes the strength of the influence: 00 means no influence, and a higher value means stronger influence. The normalization condition ensures that the total influence on each agent sums to 1. The set $N(i)$ represents the neighbors (or contacts) of agent $i$.
Let us recall some notions from graph theory [14]. A sequence in $E$ of the form $(i,i_{1})(i_{1},i_{2})\ldots(i_{m-1},j)$ is a path (of length $m$) from $i$ to $j$. The graph $G$ is said to be strongly connected if for every pair $(i,j)$ of distinct nodes in $A$, there is a path from $i$ to $j$. A graph $G$ is a clique if for every pair $(i,j)$ of distinct nodes in $A$, $(i,j)\in E$. A cycle is a path of the form $(i,i_{1})(i_{1},i_{2})\ldots(i_{m-1},i)$ with all $i_{1},\ldots i_{m-1}$ being distinct. Finally, $G$ is aperiodic if the greatest common divisor of the lengths of its cycles is one.
@@ 2.2 Silence Opinion Models
To incorporate the Spiral of Silence into the DeGroot framework, we will model the evolution of agents’ opinions alongside their decisions to remain silent about a given underlying statement or proposition. Such a proposition could include controversial statements like, for example, “AI poses a threat to humanity” or “pineapple belongs on pizza”. Therefore, the state of the agents (system) with respect to the proposition involves both the state of opinion and the state of silence.
The state of opinion of all agents is represented as a vector in $[0,1]^{n}$. If $\mathbf{B}$ is a state of opinion, then $\mathbf{B}_{i}$ denotes the opinion of agent $i$ with respect a given proposition. If $\mathbf{B}_{i}=0$ ($\mathbf{B}_{i}=1$) agent $i$ complete disagrees (agrees) with the proposition. The higher the value, the stronger the agreement.
The state of silence is represented as a vector in $\{0,1\}^{n}$. If $\mathbf{S}$ is a state of silence, $\mathbf{S}_{i}=0$ ($\mathbf{S}_{i}=1$) indicates agent $i$ is silent (is not silent) [^1] Some readers may feel that we should have used $\mathbf{S}_{i}=1$ to indicate agent $i$ is silent. Our choice, however, simplifies the update equations in the next sections.
At each discrete time unit $t\in\mathbb{N}$, every agent $i\in A$ updates their opinion and their silence state. We shall use $\mathbf{B}^{t}$ and $\mathbf{S}^{t}$ to denote the state of opinion and silence at time $t$. We can now define a general Silence DeGroot opinion model as follows.
Definition 2 (Silence Opinion Model)
@@ 2.3 Spiral of Silence Models
To build some intuition, we recall the opinion update of the DeGroot model, which can be expressed as in the following equation:
$\mathbf{B}^{t+1}_{i}=\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot(\mathbf{B}^{t}_{j}-\mathbf{B}^{t}_{i})$ (1)
for each $i\in A$, $t\in\mathbb{N}$. Thus, in the DeGroot model each agent updates their opinion by taking the weighted average of the opinion differences (disagrements) of those who influence them.
We now generalize the above DeGroot update (Eq. 1) as opinion update functions that depend not only on the current opinion state but also on the state silence and, possibly, on previous states.
@@ 2.3.1 Memoryless Update
Our first update corresponds to Option (1) mentioned earlier in Section 2: The opinions of silent contacts are ignored in the update. This can be easily realized by modifying Eq. 1 as shown in the following opinion update equation:
@@
We now define the corresponding silence update function following the Spiral of Silence Theory. First, we need some notation. Let $x,y,\tau\in[0,1]$. The $\tau$-proximity relation $x\sim_{\tau}y$ holds true iff $|x-y|\leq\tau$, i.e., if $x$ and $y$ are within a tolerance radius $\tau_{.}$ Also let $N_{i}^{t}=\{j\in N_{i}:\mathbf{S}_{j}^{t}=1\}$ be the sets of non-silent neighbours of $i$ at time $t.$ The silence update function is given as follows:
$\mathbf{S}^{t+1}_{i}=\begin{cases}1&\text{if }\left\lceil\frac{|N_{i}^{t}|}{2}\right\rceil\leq|\{j\in N_{i}^{t}\ |\ \mathbf{B}^{t}_{i}\sim_{\tau_{i}}\mathbf{B}^{t}_{j}\}|\\ 0&\text{otherwise }\end{cases}$ (3)
where $\tau_{i}\in[0,1]$ is a tolerance radius constant for agent $i$.
Intuitively, an agent $i$ considers the opinion of $j$ to be close enough to theirs if it is within their tolerance radio. The agents decide to be silent if and only if the opinions of the majority of their non-silent contacts are not close enough to theirs.
We can now define the memoryless model for the spiral of silence.
@@
Let $M=(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an SO model with $G=(A,E,I)$. Then $M$ is said to be a memoryless SO model ($\mbox{SOM}^{-}$) if for each $i\in A$ and $t\in\mathbb{N}$, $\mu_{G}(\mathbf{B}^{t},\mathbf{S}^{t},t)=(\mathbf{B}^{t+1},\mathbf{S}^{t+1})$ where $\mathbf{B}^{t+1}_{i}$ and $\mathbf{S}^{t+1}_{i}$ are determined by Eq .2 and Eq .3, respectively.
Clearly, we can recover the DeGroot update ( Eq .1), and hence the model, by setting each tolerance radius constant $\tau_{i}$ in Eq .3 to 1 and the initial state of silence $\mathbf{S}^{0}$ to the unit vector $\mathbf{1}_{n}=(1,1,\dots,1)$ of size $n$.
@@
It is worth noting that the dynamic nature of the influence graph in $\mbox{SOM}^{-}$ models sets them apart from the static influence in the DeGroot model. Silencing an agent $j$ at a given time amounts to removing all edges $(j,i)\in E$ from the graph at that moment. This allows for more complex opinion formation behaviors.
Notice also that we could normalize the sum in Eq .2 by dividing it by $\sum_{j\in N_{i}^{t}}I_{ji}$ when non-zero. While this would not impact our technical results in Section 3, it would amplify the influence of the non-silent neighbors of $i$ at time $t$, which may seem unnatural.
Instead, notice that since the right-hand side of Eq. 2 is equivalent to  $(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}$ then the influence $I_{ji}$ of a silent agent $j$ at time $t$ may be seen as increasing the weight of agent $i$’s opinion at that time. This can be interpreted as agent $i$ increasing confidence in their own opinion in the absence of external influence from agent $j$.
2.3.2 Memory-based Update
We now introduce the model corresponding to Option (2) mentioned in Section 2: If a contact $j$ is silent at time $t$, the opinion update takes into account the opinion they had the last time unit $u$ (where $u\leq t$) when they were not silent. For this to be well-defined, we assume that initially all agents are not silent; i.e., $\mathbf{S}^{0}=\mathbf{1}_{n}$.
Let $\bar{t}_{j}=\max\{u\leq t\mid\mathbf{S}^{u}_{j}=1\}$. The public state of opinion at time $t$ is a state of opinion $\text{p}\mathbf{B}^{t}$ such that $\text{p}\mathbf{B}^{t}_{j}=\mathbf{B}^{\bar{t}_{j}}_{j}$ for each $j\in A.$ The following opinion update equation captures the above intuition:
$\mathbf{B}^{t+1}_{i}=\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot(\text{p}\mathbf{B}^{t}_{j}-\mathbf{B}^{t}_{i})$ (4)
The corresponding silence update tells us that an agent $i$ becomes or remains silent at time $t+1$ precisely when the public opinion of the majority of all their contacts are not close enough to their own. More precisely:
$\mathbf{S}^{t+1}_{i}=\begin{cases}1&\text{if }\left\lceil\frac{|N_{i}|}{2}\right\rceil\leq|\{\ j\in N_{i}\ \ |\ \ \mathbf{B}^{t}_{i}\sim_{\tau_{i}}\text{p}\mathbf{B}^{t}_{j}\ \}|\\ 0&\text{otherwise }\end{cases}$ (5)
where $\tau_{i}\in[0,1]$ is a tolerance radius constant for agent $i$.
The memory-based models are defined thus:

### v2

# Jesús Aranda 11    Juan Francisco Díaz 11    David Gaona 11    Frank Valencia 2233

@@ Abstract
We generalize the classic multi-agent DeGroot framework for opinion dynamics by incorporating the Spiral of Silence theory from political science, which posits that individuals may withhold their opinions when they perceive them to be in the minority. As in the original DeGroot model, the social network is represented as a weighted directed graph encoding how agents influence one another. However, agents holding minority opinions become silent, meaning they do not express their views.
We introduce two families of models. In Silence Opinion Memoryless ($\mbox{SOM}^{-}$) models, agents update their opinions by averaging those of their non-silent neighbors. In Silence Opinion Memory-based ($\mbox{SOM}^{+}$) models, agents average the opinions of all neighbors, but for silent ones, only the most recently expressed opinion is used. We show that $\mbox{SOM}^{-}$ models guarantee consensus on clique graphs but, unlike the classic DeGroot model, not on all strongly connected aperiodic graphs. For $\mbox{SOM}^{+}$ models, even cliques may fail to reach consensus, illustrating that even minimal memory can significantly affect opinion dynamics. Finally, we validate our models through large-scale simulations on small-world networks with over two million agents. The results support the Spiral of Silence theory and reveal inherent limitations to consensus in more realistic settings.
@@ 1 Introduction
Social networks play a significant role in opinion formation often having strong impact on polarization and the ability to reach consensus. Broadly, the dynamics of opinion formation in social networks involve users expressing their opinions, being exposed to the opinions of others, and potentially adapting their own views based on these interactions. Modeling these dynamics enables us to glean insights into how opinions form and spread within social networks.
The DeGroot framework [14] is one of the most prominent formalisms for opinion formation and consensus-building in social networks. In the models of this framework, a social network is represented as a weighted directed graph, where edges denote the degree to which individuals (i.e., agents) influence one another. Each agent holds an opinion, expressed as a value in $[0,1]$, indicating their level of agreement with an underlying proposition (e.g., “AI is a threat to humanity"). Agents repeatedly update their opinions by taking the weighted average of their opinion differences with those who influence them (i.e., their contacts). There is empirical evidence validating the opinion formation through averaging of the model in controlled sociological experiments [9].
Consensus, i.e., convergence to a common opinion, is a central property in models of social learning and opinion formation [17]. In fact, difficulties in achieving consensus are a sign of a polarized society. A fundamental result for (classic) DeGroot models shows that agents converge to consensus if the influence graph is strongly connected and aperiodic. The DeGroot framework continues to be a focus of research for constructing frameworks for understanding opinion formation dynamics in social networks (e.g,. [17, 4, 3, 2, 22, 2, 13, 11, 31, 30, 15, 10]).
Nevertheless, the classic DeGroot formalism makes an assumption that could be overly constraining within social network contexts. It assumes that all agents express their opinions at each time unit. This assumption, which renders models tractable, may hold in some controlled scenarios as participants are often encouraged to express their views freely and consistently. However, in many real-world situations, some individuals may choose not to express their opinions due to personal choice or social pressure.
@@
The relevance of the Spiral of Silence has been validated in social media environments. Large-scale empirical studies show that social media platforms do not offer alternative spaces for minority opinion expression, but instead reinforce silencing behaviors [18]. Meta-analyses confirm that the link between perceived opinion support and willingness to speak out remains strong in digital contexts [21]. Recent work finds that users are particularly sensitive to perceived disagreement among their online contacts, leading to self-censorship that can spill over into offline behavior [18, 26]. During the COVID-19 pandemic, individuals frequently conformed publicly to dominant narratives while privately holding dissenting views, further illustrating the theory’s enduring relevance [6]. These findings suggest that the dynamics of opinion suppression described by The Spiral of Silence not only persist but may be exacerbated in digital social networks.
@@
In this paper, we generalize DeGroot models into a framework where agents may choose to remain silent at a given time following the Spiral of Silence. We consider two possibilities, leading to the two families of models described below.
@@
Memoryless framework $\mbox{SOM}^{-}$. In these models, silent agents are excluded from the opinion updates of the agents they would otherwise influence. Additionally, agents become silent at a given time if their views do not align with the majority of their non-silent contacts. This framework is called the silence opinion memoryless ($\mbox{SOM}^{-}$) models, as the previous opinions of silent agents are not retained. This corresponds to a social scenario in which opinions (e.g., expressed in posts) are removed once they have been accessed.
@@
Notice that ignoring silent agents at a given time unit amounts to removing certain edges from the influence graph at that time. Thus, a fundamental distinction from DeGroot models is that $\mbox{SOM}^{-}$ models exhibit dynamic influence as edges may disappear and reappear during opinion evolution.
@@
Memory-based framework $\mbox{SOM}^{+}$. In these models, agents choose to be silent if their opinion does not align with the most recent public opinions of the majority of their contacts. Unlike in $\mbox{SOM}^{+}$ models, silent agents are not excluded from the opinion updates of the agents they influence: when their current opinion is unknown, their most recent public opinion is taken into account in the update. This framework is called the silence opinion memory-based ($\mbox{SOM}^{+}$) models, as the most recent opinion of each agent is retained.
@@
A property that distinguishes $\mbox{SOM}^{+}$ models from classic DeGroot models (and $\mbox{SOM}^{-}$ models) is that the latter are Markovian processes: The next state depends on the current state but not the past states. Thus, $\mbox{SOM}^{+}$ models are history-dependent but with very limited memory; only most recent public opinions are remembered. We will show that this minimal notion of memory has an impact on consensus.
Contributions. We make the following theoretical and experimental contributions:
  1. 1.
    We generalize the DeGroot model to incorporate key aspects of the Spiral of Silence theory. To the best of our knowledge, this is the first extension of the DeGroot framework to do so.
  2. 2.
    We show that in $\mbox{SOM}^{-}$, convergence to consensus is guaranteed in clique graphs (i.e., fully connected graphs) with more than two agents. This result highlights that consensus remains possible under Spiral of Silence dynamics, even when silent agents are excluded from updates.
  3. 3.
    We prove that, unlike in the classical DeGroot model, $\mbox{SOM}^{-}$ does not guarantee consensus in all strongly connected aperiodic graphs.
  4. 4.
    We demonstrate that $\mbox{SOM}^{+}$ fails to guarantee consensus even in clique graphs. This negative result underscores that the limited memory in $\mbox{SOM}^{+}$ can significantly alter the dynamics, making consensus harder to achieve under Spiral of Silence assumptions.
  5. 5.
    We validate our models through examples and large-scale simulations involving over two million agents on randomly generated networks with small-world properties typical of real social systems. These simulations support core claims of the Spiral of Silence theory, particularly the reinforcement of dominant views in social networks. The simulation code is available at: https://github.com/DavidGaona/belief_evolution_simulator.git
@@
All in all, this paper highlights the impact of silence dynamics and memory on opinion formation and highlights the limitations of consensus in more nuanced models.
@@
The paper is organized as follows: The new silence opinion models are introduced in Section 2. The study of consensus for these models is presented in Section 3. Our large-scale simulations and case studies emerging from our models and highlighting the spiral of silence effect are presented in Section 4. The concluding remarks are given in Section 5. For the sake of space, all proofs have been moved to the Appendix.
2 Opinion Models
@@
In the DeGroot framework [14], each agent updates their opinion by taking the weighted average of the opinions of those who influence them. The models of this framework, however, do not account for the social phenomenon known as the Spiral of Silence [23], where some agents may choose to become or remain silent if their opinion does not align with the majority. As a result, their current opinion may not influence their contacts.
@@
In this section, we generalize the DeGroot framework to take into account the Spiral of Silence. If an agent $j$ decides to be silent, there are at least two natural options when updating the opinions of the agents having $j$ as a contact: (1) agent $j$ is simply ignored in the update since their current opinion is unknown (or not public), or (2) the most recent opinion when $j$ was not silent is taken into account in the update. The former corresponds to a scenario where, for privacy purposes, opinions (messages) are removed once they have been accessed. The latter represents a typical scenario in social networks where previous opinions are kept and thus continue to influence others despite the agent’s current silence.
@@
The above options lead us to the two generalizations of DeGroot models studied in this paper: the silence opinion memoryless ($\mbox{SOM}^{-}$) models, where previous opinions are forgotten, and the silence opinion memory-based (or history-dependent) ($\mbox{SOM}^{+}$) models, where previous opinions are remembered. In both, agents become silent by a majority rule for each case. Below we introduce the elements of the models.
@@ 2.1 The Influence Graph
An ($n$-agent) influence graph is a weighted directed graph $G=(A,E,I)$, where $A=\{1,\ldots,n\}$, $E\subseteq A\times A$, and $I:A\times A\rightarrow[0,1]$ a weight function such that $I(i,j)=0{\rm\ iff\ }(i,j)\not\in E$ and for each $i\in A$, $\sum_{j\in N_{i}\cup\{i\}}I(j,i)=1$ where $N_{i}=\{j\in A\setminus\{i\}:(j,i)\in E\}.$
The vertices in $A$ represent $n$ agents of a given community or network. The set of edges $E\subseteq A\times A$ represents the (direct) influence relation between these agents; i.e., $(i,j)\in E$ means that agent $i$ (directly) influences agent $j$. The value $I(i,j)$, for simplicity written $I_{ij}$, denotes the strength of the influence: 0 means no influence, and a higher value means stronger influence. The normalization condition ensures that the total influence on each agent sums to 1. The set $N_{i}$ represents the neighbors of agent $i$.
We recall some notions from graph theory [16]. A sequence in $E$ of the form $(i,i_{1})(i_{1},i_{2})\ldots(i_{m-1},j)$ is a path (of length $m$) from $i$ to $j$. The graph $G$ is said to be strongly connected if for every pair $(i,j)$ of distinct nodes in $A$, there is a path from $i$ to $j$. A graph $G$ is a clique if for every pair $(i,j)$ of distinct nodes in $A$, $(i,j)\in E$. A cycle is a path $(i,i_{1})(i_{1},i_{2})\ldots(i_{m-1},i)$ with all $i,i_{1},\ldots i_{m-1}$ being distinct. Finally, $G$ is aperiodic if the greatest common divisor of the lengths of its cycles is one.
@@ 2.2 Silence Opinion Models
To incorporate the Spiral of Silence into the DeGroot framework, we will model the evolution of agents’ opinions alongside their decisions to remain silent about a given underlying statement or proposition. Such a proposition could include controversial statements like, for example, “AI poses a threat to humanity” or “pineapple belongs on pizza”. Thus, the state of the agents (system) with respect to the proposition involves both the state of opinion and the state of silence.
The state of opinion of all agents is represented as a vector in $0,1]^{n}$. If $\mathbf{B}$ is a state of opinion, then $\mathbf{B}_{i}$ denotes the opinion of agent $i$ with respect to a given proposition. If $\mathbf{B}_{i}=0$ ($\mathbf{B}_{i}=1$) agent $i$ completely disagrees (agrees) with the proposition. The higher the value, the stronger the agreement.
The state of silence is represented as a vector in $\{0,1\}^{n}$. If $\mathbf{S}$ is a state of silence, $\mathbf{S}_{i}=1$ ($\mathbf{S}_{i}=0$) means that agent $i$ is speaking (is not speaking; i.e, agent $i$ is silent).
At each time $t\in\mathbb{N}$, every agent $i\in A$ updates their opinion and their silence state. We shall use $\mathbf{B}^{t}$ and $\mathbf{S}^{t}$ to denote the state of opinion and silence at time $t$. We can now define a general Silence DeGroot opinion model as follows.
Definition 2 (SO Models)
@@ 2.3 Spiral of Silence Models
To build intuition we recall that the opinion update from the DeGroot model states that each agent adopts the weighted average of the opinions of the agents that directly influence them. This update can be equivalently expressed in terms of opinion differences with their neighbors, as shown in the rightmost formula of the following equation:
$\mathbf{B}^{t+1}_{i}=\sum_{j\in N_{i}\cup\{i\}}I_{ji}\cdot\mathbf{B}^{t}_{j}\ =\ \mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot(\mathbf{B}^{t}_{j}-\mathbf{B}^{t}_{i})$ (1)
for each $i\in A$, $t\in\mathbb{N}$. Thus, in the DeGroot model each agent updates their opinion with the weighted average of the opinion differences with their neighbors.
We now generalize the above DeGroot update (Eq. 1) as opinion update functions that depend not only on the current state of opinion but also on the state of silence and, possibly, on previous states.
@@ 2.3.1 Memoryless Update
Our first update corresponds to the first option mentioned at the beginning of Sec. 2: The opinions of silent neighbors are ignored in the update. This can be realized by modifying Eq. 1 as shown in the following opinion update equation:
@@
We now define the corresponding silence update function following the Spiral of Silence Theory. First, we need some notation. Let $x,y,\tau\in[0,1]$. The $\tau$-proximity relation $x\sim_{\tau}y$ holds true iff $|x-y|\leq\tau$, i.e., if $x$ and $y$ are within a tolerance radius $\tau_{.}$ Also, let $N_{i}^{t}=\{j\in N_{i}:\mathbf{S}_{j}^{t}=1\}$ be the sets of non-silent neighbors of $i$ at time $t.$ The silence update function is given as follows:
[…]
For each agent $i$, the constants $\tau_{i},\mathcal{M}_{i}\in[0,1]$ represent the agent’s tolerance radius and majority threshold i.e., the minimum proportion of neighbors required for a group to be regarded as a majority by agent ii. [^1] Although the constraint $\mathcal{M}_{i}\geq 0.5$ may seem natural, we neither require nor assume it in our technical results or simulations.
Intuitively, an agent $i$ considers the opinion of $j$ to be close enough to theirs if it is within their tolerance radius $\tau_{i}$. Agent $i$ decides to speak iff at least a fraction $\mathcal{M}_{i}$, their majority threshold, of their non-silent contacts have opinions close enough to theirs.
We can now define the memoryless models for the spiral of silence.
@@
Let $M=(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an SO model with $G=(A,E,I)$. Then $M$ is said to be a SO Memoryless ($\mbox{SOM}^{-}$) model if for each $i\in A$ and $t\in\mathbb{N}$, $\mu_{G}(\mathbf{B}^{t},\mathbf{S}^{t},t)=(\mathbf{B}^{t+1},\mathbf{S}^{t+1})$ where $\mathbf{B}^{t+1}_{i}$ and $\mathbf{S}^{t+1}_{i}$ are determined by Eq. 2 and Eq. 3, respectively.
Clearly, we can recover the DeGroot update (Eq. 1) by setting each tolerance radius constant $\tau_{i}$ in Eq. 3 to 1 (or by setting each majority threshold $\mathcal{M}_{i}$ to 0) and the initial state of silence $\mathbf{S}^{0}$ to the unit vector $\mathbf{1}_{n}=(1,1,\dots,1)$ of size $n$.
@@
The dynamic nature of the influence graph in $\mbox{SOM}^{-}$ models sets them apart from the static influence in the DeGroot model. Silencing an agent $j$ at a given time amounts to removing all edges $(j,i)\in E$ from the graph at that moment. This allows for more complex opinion formation behaviors.
Furthermore, we could have normalized the sum in Eq. 2 by dividing it by $\sum_{j\in N_{i}^{t}}I_{ji}$ when this divisor is not equal to zero.  [^2] Notice we are using $N_{i}^{t}$ as index set in the summation rather than $N_{i}$. While this would not impact our technical results in Section 3, it would amplify the influence of the non-silent neighbors of $i$ at time $t$, which may seem unnatural. Instead, notice that from Eq. 2 we get:
$\mathbf{B}^{t+1}_{i}=\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot(\mathbf{B}^{t}_{j}-\mathbf{B}^{t}_{i})=(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}.$ (4)
Thus, the influence $I_{ji}$ of a silent agent $j$ at time $t$ may be seen as increasing the weight of agent $i$’s opinion at that time. This can be interpreted as agent $i$ increasing confidence in their own opinion in the absence of external influence from agent $j$.
2.3.2 Memory-based Update.
We now introduce the models corresponding to second option in the beginning of Sec. 2: If $j$ is silent at time $t$, the opinion update takes into account the opinion they had the last time unit $u$ (where $u\leq t$) when they were not silent. For this to be well-defined, we assume that initially all agents are not silent; i.e., $\mathbf{S}^{0}=\mathbf{1}_{n}$.
Let $\bar{t}_{j}=\max\{u\leq t\mid\mathbf{S}^{u}_{j}=1\}$. The public state of opinion at time $t$ is a state of opinion $\mathtt{pub}\mathbf{B}^{t}$ such that $\mathtt{pub}\mathbf{B}^{t}_{j}=\mathbf{B}^{\bar{t}_{j}}_{j}$ for each $j\in A.$ The following opinion update equation captures the above intuition:
$\mathbf{B}^{t+1}_{i}=\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot(\mathtt{pub}\mathbf{B}^{t}_{j}-\mathbf{B}^{t}_{i})$ (5)
The corresponding silence update tells us that an agent $i$ becomes or remains silent at time $t+1$ precisely when the public opinion of the majority of all their neighbors are not close enough to their own. More precisely:
[…]
where $\tau_{i},\mathcal{M}_{i}\in[0,1]$ are the tolerance radius and the majority threshold constants, respectively, for agent $i$. The memory-based models are defined thus: